OpenStudy (anonymous):

Suppose circle B, with a radius of 10 feet, has an angle ABC in standard position. Find the radian measure of the angle ABC with arc AB measuring 5pi/17 feet. Round to the nearest hundredth using 3.14 for pi. A: 9.24 B: .09 C: .92 D: None of these E: .94 Please help me! I dont understand what to do! It would be really nice if I could get the answer and how you did the problem so that way I can use this as a reference to other problems like this! thank you so much for your time and consideration!

OpenStudy (jdoe0001):

so the radius is 10 the arc's length is \(\bf \cfrac{5\pi}{17}\) so \(\bf s = r\theta \implies \theta = \cfrac{s}{r}\)

OpenStudy (anonymous):

I feel like I understand, would it be C: .92?

OpenStudy (jdoe0001):

well, the arc's length is, yes

OpenStudy (wolf1728):

I drew a diagram and it might help you understand it better